# Chunk Based Speech Pre-training with High Resolution Finite Scalar Quantization

### v1

@@ 1 Introduction
Transcribe speech into text in real time is critical for applications requiring immediate feedback, such as real-time transcription of broadcast contents, voice assistants, and simultaneous speech translation etc. It requires processing audio incrementally as it is received, rather than waiting for the entire utterance to be available. For neural network based end-to-end systems, the requirement includes two-fold. First, a speech encoder, such as causal encoder (Zhang et al., 2020b) and chunk encoder (Tsunoo et al., 2019; Shi et al., 2021), is able to process input audio cumulatively without dependency on the future input. Second, a decoder is required to decode transcription incrementally based on partial encoder outputs. Frame-Synchronous decoders (Dong et al., 2020), such as CTC (Graves et al., 2013) Transducer (Graves, 2012) CIF (Dong & Xu, 2020) and TAED (Tang et al., 2023), are capable of generating transcription based on partial encoder outputs and meet the streaming requirement naturally.
Self-supervised pre-training leverages vast amounts of unlabeled data and learn universal feature representations for downstream tasks, especially for tasks with limited supervised training data. Speech pre-training methods have emerged as the backbone of many speech processing tasks (Chung et al., 2018; Baevski et al., 2020b; Hsu et al., 2021b; Chen et al., 2021; Chiu et al., 2022; Tang et al., 2022; Baade et al., 2025). Those methods are designed for the speech applications with full utterance available, and compromises have to be made if the downstream tasks are streaming applications. Chiu et al. (2022) propose to conduct pre-training with a causal encoder, which sacrifices right context information. Fu et al. (2024) modify the encoder structure and conduct continual pre-training to adapt the model for streaming applications.
@@
In this study, we focus on building an encoder suitable for both streaming and offline modes, with a chunkwise self-supervised learning (Chunk SSL) framework as depicted in Figure 1. Chunk SSL aims to restore discrete indices of the masked frames based on unmasked frames in the last chunk and previous chunks. The discrete indices are estimated with finite scalar quantization (Mentzer et al., 2024) (FSQ) with vocabulary size up to millions (Section 3).The pre-training FSQ token has more fine-grained resolution compared with the downstream modeling units, such as phonemes or sentencepiece tokens, which usually are with vocabulary size ranging from tens to tens of thousands. We hypothesize that speech frames associated with a high resolution FSQ token are mainly mapped to an unique modeling unit in the downstream task and it makes the knowledge transfer easier from the pre-training stage to the fine-tuning stage. However, those high resolution FSQ tokens pose a great challenge for modeling and we propose to decompose a large codebook into small channel based sub-codebooks to alleviate the memory and computation cost during pre-training. Details could be found in Section 3.
@@ 2 Copy and Append Data Augmentation
The naive Chunk SSL algorithm, which recovers masked frame indices in the right most chunk based on the unmasked frames in the same chunk and preceding chunks, is executed chunk by chunk in a sequential order instead of computing all chunks from one utterance in a parallel fashion. Inspired by the implementation of the chunk encoder with a look-ahead chunk for the streaming speech translation (Liu et al., 2021), we propose a copy and append data augmentation for the Chunk SSL, which reorganizes input data and changes the augmented data computation, but it is still strictly equivalent to the computation in the naive Chunk SSL algorithm.
@@
A speech encoder, such as Transformer (Vaswani et al., 2017) and Conformer (Gulati et al., 2020), contains two types of modules. One is the intra-frame module where the computation is based on one frame and no location information or other frames are required, for example, LayerNorm and feedforward layers. The other type is the inter-frame module where the computation is related to the input location and requires information from neighbouring frames. The inter-frame modules include self-attention and convolutional layers. An augmented utterance aforementioned, can be processed as a normal utterance in the intra-frame modules and no modification is required. On the other hand, the position information is invalid among extended chunk frames, since they are copied and appended to the end of utterance. Modifications are introduced for two inter-frame modules: self-attention and convolutional layers, to handle inconsistent position information in augmented utterances as described in the following subsections.
@@ 2.1 CADA self-attention module
where is a scaling factor, are transform matrices, and is the dimension of model embedding111Single head attention is discussed for simplicity.. is a masking matrix with value 1 or 0 for each component and means will not contribute the computation of attention weights and the corresponding attention output of . The self-attention module leverages the attention mask to control information exchange within frames from the same sequence. For example, an encoder usually sets all for full information access.
@@ 2.2 CADA convolutional module
A convolutional layer has a set of filters which slide across input features within the same utterance and extract localized representations. A convolutional layer assumes the input frames organized as consecutively, which no longer holds true for extended chunk frames in a CADA utterance. This problem can be resolved via a chunk based convolutional computation (Li et al., 2023). Without loss of generality, we assume the convolution module is a depth-wise convolution layer with left and right context frames. The modified computation includes three steps. First, we separate the CADA utterance into chunks and pair a base chunk and its extended chunk together, for example “B1” and “E1” in Figure 1; then the last frames from the chunk preceding the base chunk are appended to the left as left context. If the base chunk is the first chunk, zero padding frames are attached. Similarly, zero padding frames are appended to the right and it leads to a new concatenated subsequence with length . Second, the concatenated subsequence is fed to the convolution module and obtains output subsequence with length . and context frames are absorbed during convolution computation. The first output frames correspond to the base chunk and the second frames are for the extended chunk. Finally, those output chunks are placed back to their original positions in the augmented utterance as the outputs from the CADA convolutional module. 2(b) demonstrates three steps to conduct convolution computation on an augmented utterance with base sequence length 12 and chunk size 4.
@@ 2.3 Dynamic chunk based pre-training
Dynamic chunk training (Zhang et al., 2020a; Weninger et al., 2022; Li et al., 2023) is a useful fine-tuning approach for the chunk based ASR encoder. During training, the chunk duration varied from hundreds of million seconds to a few seconds is randomly chosen for every model update. A model built with dynamic chunk training is suitable for both streaming and offline modes. We extend dynamic chunk training for the Chunk SSL pre-training. Chunk duration is chosen from 6 durations ( million seconds) randomly for every model update.
@@ 3.1 Quantization module
As shown in Figure 1, we leverage finite scalar quantization (Mentzer et al., 2024) to quantize input frames and obtain their discrete indices. A FSQ module is built prior to the Chunk SSL pre-training. In the quantization module, an utterance based channel mean and variance normalization is first applied to input frames to obtain , then it is processed by a FSQ encoder and projected into a low-dimension space with size , where . The output of each channel for input frame is rounded to an integer
@@ 4.1 Data
Pre-training: There are two pre-training tasks in this work, i.e., FSQ SSL training and Chunk SSL training. We first build FSQ module to discretize the input speech feature and then pre-train speech encoder with Chunk SSL to generate contextual representation. Libri-light (Kahn et al., 2019), which includes 60k hours of unlabelled English speech data, is used in both pre-training tasks.
Fine-Tuning: For speech recognition task, the pre-trained models are finetuned and evaluated on Librispeech (Panayotov et al., 2015) dataset. We use two dev sets for development and report all results from dev and test sets. For speech translation task, experiments are conducted on two MuST-C (Gangi et al., 2019) language pairs: English to German (ENDE) and English to Spanish (ENES). Sequence level knowledge distillation (Kim & Rush, 2016) is applied to boost the speech translation quality. The models are developed on the dev set, and the final results are reported on the tst-COMMON set.
@@ 4.2 Model configures
The speech encoder is a Conformer (Gulati et al., 2020) based chunk encoder. The encoder is equipped with a relative positional embedding (Shaw et al., 2018) and starts with a stacking layer to down-sample the input features by four times. Two model configurations: base and large, have been explored. The base encoder has 12 Conformer layers, input embedding size of 512, 8 attention heads, feedforward layer dimension 2048 and convolutional module kernel size 31. The large encoder has 24 Conformer layers, input embedding size of 768, 16 attention heads, feedforward layer dimension 3072 and convolutional module kernel size 5. Transducer is adopted in the fine-tuning experiments. The predictor module is with one Transformer layer (Vaswani et al., 2017) for speech recognition tasks and two layers for speech translation tasks. The input embedding and feedforward layer dimension are the same as ones in the encoder setting if not mentioned specifically. The joint module is a feedforward layer as (Zhang et al., 2020b) with embedding dimension 1024.
Input speech is represented as 80-dimensional log mel-filterbank coefficients computed every 10ms with a 25ms window. Global channel mean and variance normalization is applied so the trained model could be used for both streaming and offline scenarios. We set the maximum utterance duration to 75 seconds and minimum duration to 5 seconds in pre-training. During fine-tuning, a look-ahead chunk (Shi et al., 2021; Liu et al., 2021; Tang et al., 2023) is utilized and dynamic chunk training is employed by default. We alternate between offline training with infinite chunk size, and streaming training with chunk duration sampled from [160, 320, 640, 960, 1280, 1600] ms randomly within each epoch. More details about optimization, such as batch sizes and training schedulers, for different experiments are presented in Appendix A. The target labels are encoded with SentencePiece (Kudo & Richardson, 2018). For both speech recognition and speech translation tasks, the vocabulary is an unigram model with size 1024 and full character coverage on the corresponding training text data.
@@ 4.3 FSQ module
The FSQ module is built with the ResNet based encoder and decoder (Langman et al., 2024). There are 12 layers for both encoder and encoder with embedding size 512. The frontend processing is the same as the one described in Section 4.2. The mel-filterbank feature is with a 25ms window and 10ms shift. 4 mel-filterbank stacked features are used as input for the FSQ encoder and the reconstructed target for the FSQ decoder. The default encoder output dimension is 12 with number of levels per channel . It is equivalent to a codebook with vocabulary size 6,834,375.
@@ 5.1 Main results
| wav2vec2 base (Baevski et al., 2020b) | 95 | 3.2 | 8.9 | 3.4 | 8.5 | 6.0 |
| wav2vec2 large (Baevski et al., 2020b) | 317 | 2.1 | 4.5 | 2.2 | 4.5 | 3.3 |
| w2v-Conformer XL (Zhang et al., 2020c) | 600 | 1.7 | 3.5 | 1.7 | 3.5 | 2.6 |
| BEST-RQ (Chiu et al., 2022) | 600 | 1.5 | 2.8 | 1.6 | 2.9 | 2.2 |
@@
| Conformer (Chiu et al., 2022) | 100 | 4.1 | 10.3 | 4.5 | 9.8 | 7.2 |
| Conformer (Chiu et al., 2022) | 600 | 3.9 | 9.8 | 4.4 | 9.4 | 6.9 |
| wav2vec2 (Chiu et al., 2022) | 600 | 2.7 | 8.0 | 2.9 | 7.9 | 5.4 |
| w2v-BERT (Chiu et al., 2022) | 600 | 2.7 | 8.4 | 3.0 | 8.1 | 5.6 |
| BEST-RQ (Chiu et al., 2022) | 600 | 2.5 | 6.9 | 2.8 | 6.6 | 4.7 |
@@ 5.2 Impact of FSQ codebook sizes
In order to verify our a high resolution FSQ codebook hypothesis discussed in Section 3, we study FSQ tokenizers with different codebook sizes. Two aspects are examined: first the agreement among FSQ tokens and phonemes, and second WERs for ASR models initialized with different FSQ tokenizers. For the agreement among FSQ tokens and phonemes, we compare the overlap of features associated with different tokens and phonemes. It can be measured by phone purity and phone-normalized mutual information (PNMI) (Hsu et al., 2021a). Phone purity denotes the conditional probability of a phoneme given a FSQ index while PNMI measures the percentage of uncertainty about a phoneme label eliminated after observing a specific FSQ token. We generate phoneme based alignment with Montreal Aligner (McAuliffe et al., 2017)333https://github.com/MontrealCorpusTools/mfa-models/releases/download/acoustic-english_us_arpa-v3.0.0/english_us_arpa.zip on two Librispeech dev sets.

### v2

@@ 1 Introduction
Transcribe speech into text in real time is critical for applications requiring immediate feedback, such as real-time transcription of broadcast contents, voice assistants, and simultaneous speech translation etc. It requires processing audio incrementally as it is received, rather than waiting for the entire utterance to be available. For neural network based end-to-end systems, the requirement includes two-fold. First, a speech encoder, such as causal encoder (Zhang2020TransformerTA) and chunk encoder (Tsunoo2019TransformerAW; Shi2020EmformerEM), is able to process input audio cumulatively without dependency on the future input. Second, a decoder is required to decode transcription incrementally based on partial encoder outputs. Frame-Synchronous decoders (Dong2020ACO), such as CTC (Graves2013SpeechRW) Transducer (Graves2012SequenceTW) CIF (Dong2020CIFCI) and TAED (Tang2023HybridTA), are capable of generating transcription based on partial encoder outputs and meet the streaming requirement naturally.
Self-supervised pre-training leverages vast amounts of unlabeled data and learn universal feature representations for downstream tasks, especially for tasks with limited supervised training data. Speech pre-training methods have emerged as the backbone of many speech processing tasks (Chung2018UnsupervisedCA; Baevski2020wav2vecV2; Hsu2021HubertHM; Chen2021WavLMLS; Chiu2022SelfsupervisedLW; Tang2022UnifiedSP; Baade2024SyllableLMLC). Those methods are designed for the speech applications with full utterance available, and compromises have to be made if the downstream tasks are streaming applications. Chiu2022SelfsupervisedLW propose to conduct pre-training with a causal encoder, which sacrifices right context information. Fu2024wav2vecSAP modify the encoder structure and conduct continual pre-training to adapt the model for streaming applications.
@@
In this study, we focus on building an encoder suitable for both streaming and offline modes, with a chunkwise self-supervised learning (Chunk SSL) framework as depicted in Figure 1. Chunk SSL aims to restore discrete indices of the masked frames based on unmasked frames in the last chunk and previous chunks. The discrete indices are estimated with finite scalar quantization (Mentzer2023FiniteSQ) (FSQ) with vocabulary size up to millions (Section 3).The pre-training FSQ token has more fine-grained resolution compared with the downstream modeling units, such as phonemes or sentencepiece tokens, which usually are with vocabulary size ranging from tens to tens of thousands. We hypothesize that speech frames associated with a high resolution FSQ token are mainly mapped to an unique modeling unit in the downstream task and it makes the knowledge transfer easier from the pre-training stage to the fine-tuning stage. However, those high resolution FSQ tokens pose a great challenge for modeling and we propose to decompose a large codebook into small channel based sub-codebooks to alleviate the memory and computation cost during pre-training. Details could be found in Section 3.
@@ 2 Copy and Append Data Augmentation
The naive Chunk SSL algorithm, which recovers masked frame indices in the right most chunk based on the unmasked frames in the same chunk and preceding chunks, is executed chunk by chunk in a sequential order instead of computing all chunks from one utterance in a parallel fashion. Inspired by the implementation of the chunk encoder with a look-ahead chunk for the streaming speech translation (liu2021cross), we propose a copy and append data augmentation for the Chunk SSL, which reorganizes input data and changes the augmented data computation, but it is still strictly equivalent to the computation in the naive Chunk SSL algorithm.
@@
A speech encoder, such as Transformer (Vaswani2017AttentionIA) and Conformer (Gulati2020ConformerCT), contains two types of modules. One is the intra-frame module where the computation is based on one frame and no location information or other frames are required, for example, LayerNorm and feedforward layers. The other type is the inter-frame module where the computation is related to the input location and requires information from neighboring frames. The inter-frame modules include self-attention and convolutional layers. An augmented utterance aforementioned, can be processed as a normal utterance in the intra-frame modules and no modification is required. On the other hand, the position information is invalid among extended chunk frames, since they are copied and appended to the end of utterance. Modifications are introduced for two inter-frame modules: self-attention and convolutional layers, to handle inconsistent position information in augmented utterances as described in the following subsections.
The psuedo code for the CADA modified Conformer encoder is presented in Algorithm 1, where is number of encoder layers, CONCAT, LN, FFN and Conv are concatenation operator, layernorm, feedforward and convolutionaly modules respectively. The details of CADA self-attention and convolution modules are described in the following subsection.
@@ 2.1 CADA self-attention module
where is a scaling factor, is a big number (1e6), are transform matrices, and is the dimension of model embedding111Single head attention is discussed for simplicity.. is a masking matrix with value 1 or 0 for each component and means will not contribute the computation of attention weights and the corresponding attention output of . The self-attention module leverages the attention mask to control information exchange within frames from the same sequence. For example, an encoder usually sets all for full information access.
@@ 2.2 CADA convolutional module
A convolutional layer has a set of filters which slide across input features within the same utterance and extract localized representations. A convolutional layer assumes the input frames organized as consecutively, which no longer holds true for extended chunk frames in a CADA utterance. This problem can be resolved via a chunk based convolutional computation (Li2023DynamicCC). Without loss of generality, we assume the convolution module is a depth-wise convolution layer with left and right context frames. The modified computation includes three steps. First, we separate the CADA utterance into chunks and pair a base chunk and its extended chunk together, for example “B1” and “E1” in Figure 1; then the last frames from the chunk preceding the base chunk are appended to the left as left context. If the base chunk is the first chunk, zero padding frames are attached. Similarly, zero padding frames are appended to the right and it leads to a new concatenated subsequence with length . Second, the concatenated subsequence is fed to the convolution module and obtains output subsequence with length . and context frames are absorbed during convolution computation. The first output frames correspond to the base chunk and the second frames are for the extended chunk. Finally, those output chunks are placed back to their original positions in the augmented utterance as the outputs from the CADA convolutional module. 2(b) demonstrates three steps to conduct convolution computation on an augmented utterance with base sequence length 12 and chunk size 4.
@@ 2.3 Dynamic chunk based pre-training
Dynamic chunk training (Zhang2020UnifiedSA; Weninger2022ConformerWD; Li2023DynamicCC) is a useful fine-tuning approach for the chunk based ASR encoder. During training, the chunk duration varied from hundreds of million seconds to a few seconds is randomly chosen for every model update. A model built with dynamic chunk training is suitable for both streaming and offline modes. We extend dynamic chunk training for the Chunk SSL pre-training. Chunk duration is chosen from 6 durations ( million seconds) randomly for every model update.
@@ 3.1 Quantization module
As shown in Figure 1, we leverage finite scalar quantization (Mentzer2023FiniteSQ) to quantize input frames and obtain their discrete indices. A FSQ module is built prior to the Chunk SSL pre-training. In the quantization module, an utterance based channel mean and variance normalization is first applied to input frames to obtain , then it is processed by a FSQ encoder and projected into a low-dimension space with size , where . The output of each channel for input frame is rounded to an integer
@@ 4.1 Data
Pre-training: There are two pre-training tasks in this work, i.e., FSQ SSL training and Chunk SSL training. We first build FSQ module to discretize the input speech feature and then pre-train speech encoder with Chunk SSL to generate contextual representation. Libri-light (Kahn2019LibriLightAB), which includes 60k hours of unlabelled English speech data, is used in both pre-training tasks.
Fine-Tuning: For speech recognition task, the pre-trained models are finetuned and evaluated on Librispeech (Panayotov2015LibrispeechAA) dataset. We use two dev sets for development and report all results from dev and test sets. For speech translation task, experiments are conducted on two MuST-C (Gangi2019MuSTCAM) language pairs: English to German (ENDE) and English to Spanish (ENES). Sequence level knowledge distillation (Kim2016SequenceLevelKD) is applied to boost the speech translation quality. The models are developed on the dev set, and the final results are reported on the tst-COMMON set.
@@ 4.2 Model configures
The speech encoder is a Conformer (Gulati2020ConformerCT) based chunk encoder. The encoder is equipped with a relative positional embedding (Shaw2018SelfAttentionWR) and starts with a stacking layer to down-sample the input features by four times. Two model configurations: base and large, have been explored. The base encoder has 12 Conformer layers, input embedding size of 512, 8 attention heads, feedforward layer dimension 2048 and convolutional module kernel size 31. The large encoder has 24 Conformer layers, input embedding size of 768, 16 attention heads, feedforward layer dimension 3072 and convolutional module kernel size 5. Transducer is adopted in the fine-tuning experiments. The predictor module is with one Transformer layer (Vaswani2017AttentionIA) for speech recognition tasks and two layers for speech translation tasks. The input embedding and feedforward layer dimension are the same as ones in the encoder setting if not mentioned specifically. The joint module is a feedforward layer as (Zhang2020TransformerTA) with embedding dimension 1024.
Input speech is represented as 80-dimensional log mel-filterbank coefficients computed every 10ms with a 25ms window. Global channel mean and variance normalization is applied so the trained model could be used for both streaming and offline scenarios. We set the maximum utterance duration to 75 seconds and minimum duration to 5 seconds in pre-training. During fine-tuning, a look-ahead chunk (Shi2020EmformerEM; liu2021cross; Tang2023HybridTA) is utilized and dynamic chunk training is employed by default. We alternate between offline training with infinite chunk size, and streaming training with chunk duration sampled from [160, 320, 640, 960, 1280, 1600] ms randomly within each epoch. More details about optimization, such as batch sizes and training schedulers, for different experiments are presented in Appendix A. The target labels are encoded with SentencePiece (Kudo2018SentencePieceAS). For both speech recognition and speech translation tasks, the vocabulary is an unigram model with size 1024 and full character coverage on the corresponding training text data.
@@ 4.3 FSQ module
The FSQ module is built with the ResNet based encoder and decoder (Langman2024SpectralCI). There are 12 layers for both encoder and encoder with embedding size 512. The frontend processing is the same as the one described in Section 4.2. The mel-filterbank feature is with a 25ms window and 10ms shift. 4 mel-filterbank stacked features are used as input for the FSQ encoder and the reconstructed target for the FSQ decoder. The default encoder output dimension is 12 with number of levels per channel . It is equivalent to a codebook with vocabulary size 6,834,375.
@@ 5.1 Main results
| wav2vec2 base (Baevski2020wav2vecV2) | 95 | 3.2 | 8.9 | 3.4 | 8.5 | 6.0 |
| wav2vec2 large (Baevski2020wav2vecV2) | 317 | 2.1 | 4.5 | 2.2 | 4.5 | 3.3 |
| w2v-Conformer XL (Zhang2020PushingTL) | 600 | 1.7 | 3.5 | 1.7 | 3.5 | 2.6 |
| BEST-RQ (Chiu2022SelfsupervisedLW) | 600 | 1.5 | 2.8 | 1.6 | 2.9 | 2.2 |
@@
| Conformer (Chiu2022SelfsupervisedLW) | 100 | 4.1 | 10.3 | 4.5 | 9.8 | 7.2 |
| Conformer (Chiu2022SelfsupervisedLW) | 600 | 3.9 | 9.8 | 4.4 | 9.4 | 6.9 |
| wav2vec2 (Chiu2022SelfsupervisedLW) | 600 | 2.7 | 8.0 | 2.9 | 7.9 | 5.4 |
| w2v-BERT (Chiu2022SelfsupervisedLW) | 600 | 2.7 | 8.4 | 3.0 | 8.1 | 5.6 |
| BEST-RQ (Chiu2022SelfsupervisedLW) | 600 | 2.5 | 6.9 | 2.8 | 6.6 | 4.7 |
@@ 5.2 Impact of FSQ codebook sizes
In order to verify our a high resolution FSQ codebook hypothesis discussed in Section 3, we study FSQ tokenizers with different codebook sizes. Two aspects are examined: first the agreement among FSQ tokens and phonemes, and second WERs for ASR models initialized with different FSQ tokenizers. For the agreement among FSQ tokens and phonemes, we compare the overlap of features associated with different tokens and phonemes. It can be measured by phone purity and phone-normalized mutual information (PNMI) (Hsu2021HubertSS). Phone purity denotes the conditional probability of a phoneme given a FSQ index while PNMI measures the percentage of uncertainty about a phoneme label eliminated after observing a specific FSQ token. We generate phoneme based alignment with Montreal Aligner (mcauliffe17_interspeech)333https://github.com/MontrealCorpusTools/mfa-models/releases/download/acoustic-english_us_arpa-v3.0.0/english_us_arpa.zip on two Librispeech dev sets.
@@
5.3 Comparison FSQ v.s. BEST-RQ
We compare the speech recognition performance with different initialization approaches: 1) random initialization (“Random”), 2) Chunk SSL model with discretized token created by BEST-RQ (Chiu2022SelfsupervisedLW) (“Chunk SSL + BEST-RQ”), 3) Chunk SSL model with FSQ codebook size 50,625 (“Chunk SSL + FSQ (50k)”) and 4) Chunk SSL model with FSQ codebook size 6,834,375 (“Chunk SSL + FSQ (6m)”). The results are listed in Table 5. The results show FSQ with a moderate codebook size achieves similar results as BEST-RQ, and FSQ could further improve the downstream performance by adopting a large codebook with size around 6 millions.
| Model | Mode | dev-clean | dev-other | test-clean | test-other | ave. |
| Random | Offline | 2.3 | 5.8 | 2.4 | 5.8 | 4.1 |
| Chunk SSL + BEST-RQ | Offline | 2.1 | 5.3 | 2.2 | 5.3 | 3.7 |
| Chunk SSL + FSQ (50k) | Offline | 2.1 | 5.3 | 2.2 | 5.1 | 3.7 |
| Chunk SSL + FSQ (6m) | Offline | 2.0 | 5.0 | 2.1 | 4.9 | 3.5 |
| Random | Streaming | 2.6 | 7.4 | 2.8 | 7.2 | 5.0 |
| Chunk SSL + BEST-RQ | Streaming | 2.4 | 6.8 | 2.7 | 6.8 | 4.6 |
| Chunk SSL + FSQ (50k) | Streaming | 2.4 | 6.9 | 2.5 | 6.7 | 4.6 |
| Chunk SSL + FSQ (6m) | Streaming | 2.3 | 6.6 | 2.5 | 6.4 | 4.5 |